# Study of hybrid stars with nonstrange quark matter cores

### v1

@@ Abstract
In this work, we use a modified 2-flavor Nambu-Jona-Lasinio model to study hybrid stars with a hypothesis that the quark matter may not be strange [Phys. Rev. Lett. 120, 222001 (2018)]. To obtain hybrid equation of states, the Maxwell construction is used to describe the first-order confinement-deconfinement phase transition in hybrid stars. With recent measurements on neutron star mass, radius, and tidal deformability, the hybrid equation of states are constrained. The result suggests that pure nonstrange quark matter cores can exist in hybrid stars, possessing 0.014-0.026 solar mass, and the binary neutron stars in GW170817 may be hadron stars.
@@ I INTRODUCTION
To describe the hadronic matter in hybrid stars, the Hartree-Fock-Bogoliubov (HFB) mass model HFB-21 with unconventional Skyrme forces called BSK21 Goriely et al. (2010); Chamel et al. (2011); Fantina, A. F. et al. (2013); Giliberti and Cambiotti (2022) is a good approach, and the corresponding hadronic EOS obtained in this framework is consistent with current constraints on neutron star masses, radii, and the tidal deformability from GW170817 Fantina, A. F. et al. (2022). However, to describe the quark matter in hybrid stars, the lattice QCD is confronted with difficulties at low-temperature and high-density regions because of the “sign problem”, thus we need to use effective models, such as the Nambu-Jona-Lasinio (NJL) model Klevansky (1992); Buballa (2005); Li et al. (2019a); Liu et al. (2021); Zhang et al. (2022); Huang and Zhuang (2023), which manifests the mechanism of spontaneous breaking of chiral symmetry. In this work, inspired by the operator product expansion (OPE) method, we consider to modify the four-fermion interaction coupling constant $G$ in the NJL model as $G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$ Jiang et al. (2012); Cui et al. (2013, 2014); Shi et al. (2016); Wang et al. (2016); Fan et al. (2017, 2019); Li et al. (2018a), where $G_{2}\langle\bar{\psi}\psi\rangle$ represents the feedback of the quark propagator.
@@
As for the hadron-quark phase transition, the most widely used approach is the Maxwell construction Endo et al. (2006); Hempel et al. (2009); Yasutake et al. (2009), assuming that the first-order phase transition occurs Glendenning (1992); Bhattacharyya et al. (2010) and stable quark matter cores exist in hybrid stars. However, many studies showed that hybrid stars are unstable against oscillations in this case, because star masses decrease with the increase of the central density, thus quark matter cores may not exist in neutron-star interiors Özel (2006); Hoyos et al. (2016); Qin et al. (2023). In Ref. Özel (2006) and Ref. Hoyos et al. (2016), the theoretical modeling of bursting neutron-star spectra and top-down holographic model for strongly interacting quark matter were employed, respectively, to demonstrate that the two-solar-mass (2 $M_{\odot}$) neutron star has ruled out soft EOSs of neutron-star matter, and no quark matter exists in massive neutron stars. Recently, it has been argued that as the density increases, the boundaries of hadrons disappear gradually and the corresponding phase transition is a crossover Baym et al. (2018). According to this assumption, the three-window modeling Masuda et al. (2013a, b) in the crossover region was proposed. Many studies has constructed hybrid EOSs in this scheme, and the corresponding maximum masses of hybrid stars are compatible with 2 $M_{\odot}$ojo et al. (2015); Li et al. (2017, 2018a, 2018b, 2022b); Qin et al. (2023).
@@
In addition to theoretical studies of hadron-quark phase transitions and hybrid EOSs, astronomical observations of neutron star masses, radii, and tidal deformability have also placed constraints on numerous EOSs. Some massive neutron star observations such as PSR J0348+0432 Antoniadis et al. (2013) and PSR J0740+6620 Cromartie et al. (2020) require EOSs should not be too soft, but the tidal deformability constrained in BNS merger event GW170817 indicates the EOSs should not be too stiff Abbott and et al. (2017a, 2018). Recently, the joint mass-radius observations of neutron stars from NASA’s Neutron Star Interior Composition Explorer (NICER) missions have also imposed some constraints on these EOSs Riley et al. (2019); Miller et al. (2019); Riley et al. (2021); Miller et al. (2021). In Refs. Bauswein et al. (2019); Miao et al. (2020), the authors claim that the gravitational-wave (GW) emission of GW170817 supports a first-order hadron-quark phase transition at supranuclear densities.
@@
This paper is organized as follows. In Sec.II, the modified NJL model for nonstrange quark matter is briefly introduced, and the EOSs of quark matter are derived. In Sec.III, the Maxwell construction is used to get hybrid EOSs. With recent astronomical observations of neutron star mass, radius, and tidal deformability, we constrain the hybrid EOSs. For comparison, the mass-radius relations and tidal deformability results of hybrid stars from six representative hybrid EOSs are presented. Finally, a brief summary is given in Sec.IV.
@@ III Structure of hybrid stars
Once the EOS is determined, we can solve the Tolman-Oppenheimer-Volkoff (TOV) equation numerically to get the mass-radius ($M-R$) and mass-central energy density ($M-\epsilon_{\rm c}$) relations. In Fig. 8 and Fig. 9, to study the influence of $G_{1}$ and $B^{1/4}$ on the $M-R$ and $M-\epsilon_{\rm c}$ relations, the corresponding results for (a) $B^{1/4}=153$ MeV and (b) $G_{1}=1.935$ GeV-2 are shown, respectively. In Fig. 8, the most massive quark star with $G_{1}=1.935,1.950,1.972$ GeV-2 ($B^{1/4}=150,153,161$ MeV) is about 1.52 $M_{\odot}$ (1.57 $M_{\odot}$), not satisfying the 2 $M_{\odot}$ constraint. In addition, the quark stars described by the modified NJL model in this work cannot fulfill the recent $M-R$ constraint from the NICER mission (PSR J0740+6620 Miller et al. (2021) and PSR J0030+0451 Miller et al. (2019)). However, the hybrid EOSs obtained with the Maxwell construction approach in this work can produce hybrid stars in consistent with these astronomical observations, although their quark matter cores are relatively small (about 0.02 $M_{\odot}$).
Figure 8: The $M-R$ relations of quark stars and hybrid stars for (a) $B^{1/4}=153$ MeV and (b) $G_{1}=1.935$ GeV-2. The $M-R$ relation of the hadron star described by BSK21 is shown by the red line. QS1935, QS1950, QS1972 (HS1935, HS1950, HS1972) refer to quark stars (hybrid stars) with $G_{1}=1.935,1.950,1.972$ GeV-2, respectively, and QS150, QS153, QS161 (HS150, HS153, HS161) refer to quark stars (hybrid stars) with $B^{1/4}=150,153,161$ MeV, respectively. The constraints from the NICER mission (PSR J0740+6620 Miller et al. (2021) and PSR J0030+0451 Miller et al. (2019)) are shown as dark (light) colored areas, referring to the $68\%$ ($95\%$) confidence levels.
@@ IV Summary and discussion
When the current quark mass is chosen as $m=3.5$ MeV, which is estimated by the particle data group, the modification of the coupling constant $G$ in the normal NJL model is helpful, because it is not only consistent with the QCD requirement in essence, but also in agreement with the lattice simulation results of $T_{\rm c}$. In the 2-flavor case when $T_{\rm c}=173\pm 8$ MeV, our calculations indicate that the parameter space of $G_{1}$ is limited to a range of 1.935-1.972 GeV-2. For hybrid EOSs, the influence of $G_{1}$ is very small, but the influence of $B^{1/4}$ is obvious.
Considering astronomical observations and the stability of hybrid stars, the parameter $B^{1/4}$ is constrained to be 150-161, 149-158, 148-157 MeV when $G_{1}=1.935,1.950,1.972$ GeV-2, respectively. The quark EOSs constructed with the modified NJL model in this work is soft, and thus cannot satisfy the 2 $M_{\odot}$ constraint of neutron stars and the mass-radius constraints from NICER missions. However, the hybrid EOSs with first-order hadron-quark transitions are still in agreement with these constraints, and pure nonstrange quark matter cores can exist in hybrid stars, possessing a mass of 0.014-0.026 $M_{\odot}$. According to the hybrid EOSs constrained in this work, the BNS in GW170817 whose masses are estimated to be 1.17-1.36 and 1.36-1.60 $M_{\odot}$ Abbott and et al. (2017a), respectively, may be hadron stars.
@@ Acknowledgements.
This work is supported in part by the national Key Program for Science and Technology Research Development (2023YFB3002500), the National Natural Science Foundation of China (under Grants No. 12005192, No. 12075213, and No. 12233002), the Project funded by China Postdoctoral Science Foundation (No. 2020M672255, No. 2020TQ0287), National SKA Program of China No. 2020SKA0120300, the National Key R&\&&D Program of China (2021YFA0718500), the Natural Science Foundation of Henan Province of China (No. 242300421375), the start-up funding from Zhengzhou University. Y.F.H also acknowledges the support from the Xinjiang Tianchi Program.

### v2

@@ Abstract
In this work, under the hypothesis that quark matter may not be strange [Phys. Rev. Lett. 120, 222001 (2018)], we adopt a modification of the coupling constant of the four-quark scalar interaction $G\rightarrow G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$ in the 2-flavor Nambu-Jona-Lasinio model to study nonstrange hybrid stars. According to lattice QCD simulation results of the critical temperature at zero chemical potential, $G_{1}$ and $G_{2}$ are constrained as $G_{1}\in(1.935,1.972)$ GeV-2, and $G_{2}\in(-1.582,-0.743)$ GeV-5, respectively. To obtain hybrid equation of states, the Maxwell construction is used to describe the first-order confinement-deconfinement phase transition in hybrid stars. With recent measurements on neutron star mass, radius, and tidal deformability, the hybrid equation of states are constrained. The result suggests that pure nonstrange quark matter cores can exist in hybrid stars, possessing 0.014-0.026 solar mass with the bag constant $B^{1/4}$ in a range of 148-161 MeV. It is argued that the binary neutron stars in GW170817 should be hadron stars.
@@ I INTRODUCTION
As we know, the results of different effective models can be quantitatively or even qualitatively different. Even for the same model, if different modifications are taken into account, the results can also be different. For example, Fig. 10 of Ref. Özel and Freire (2016) shows that the EOSs given by different effective models are different from each other, and the corresponding mass-radius ($M-R$) relations of neutron stars are also different. Thus there is not a definite answer to the EOS of dense strongly interacting matter at zero temperature at present.
@@
To describe the hadronic matter in hybrid stars, the Hartree-Fock-Bogoliubov (HFB) mass model HFB-21 with unconventional Skyrme forces called BSK21 Goriely et al. (2010); Chamel et al. (2011); Fantina, A. F. et al. (2013); Giliberti and Cambiotti (2022) is a good approach, and the corresponding hadronic EOS obtained in this framework is consistent with current constraints on neutron star masses, radii, and the tidal deformability from GW170817 Fantina, A. F. et al. (2022). However, to describe the quark matter in hybrid stars, the lattice QCD is confronted with difficulties at low-temperature and high-density regions because of the “sign problem”, thus we need to use effective models, such as the Nambu-Jona-Lasinio (NJL) model Klevansky (1992); Buballa (2005); Li et al. (2019a); Liu et al. (2021); Zhang et al. (2022); Huang and Zhuang (2023), which manifests the spontaneous breaking of chiral symmetry.
In the framework of the NJL-type model, many studies focused on hybrid stars Ayriyan et al. (2021); Blaschke et al. (2020); Alvarez-Castillo et al. (2016); Pfaff et al. (2022), aiming to explain the observed two-solar-mass (2 $M_{\odot}$) compact stars. In Refs. Ayriyan et al. (2021); Blaschke et al. (2020); Alvarez-Castillo et al. (2016), the 2-flavor NJL-type model was used to consider the scalar quark-antiquark interaction, anti-triplet scalar diquark interactions and vector quark-antiquark interactions Ayriyan et al. (2021), a chemical potential dependence of the vector mean-field coupling $\eta(\mu)$ and a chemical potential-dependent bag constant $B(\mu)$ Blaschke et al. (2020), multiquark (4- and 8-quark) interactions Alvarez-Castillo et al. (2016), respectively. In Ref. Pfaff et al. (2022), the 3-flavor SU(3) NJL model was adopted with the four-quark scalar, vector-isoscalar and vector-isovector interactions as well as the ’t Hooft interaction. Different from the above studies, here we adopt a modification of the coupling constant of four-quark scalar interactions as $G\rightarrow G_{1}+G_{2}\langle\bar{\psi}\psi\rangle$ in this work, which can be regarded as a representation of an effective gluon propagator (See Sec. II for specific analysis).
@@
As for the hadron-quark phase transition, the most widely used approach is the Maxwell construction Endo et al. (2006); Hempel et al. (2009); Yasutake et al. (2009), assuming that the first-order phase transition occurs Glendenning (1992); Bhattacharyya et al. (2010) and stable quark matter cores exist in hybrid stars. However, many studies showed that hybrid stars are unstable against oscillations in this case, because star masses decrease with the increase of the central density, thus quark matter cores may not exist in neutron-star interiors Özel (2006); Hoyos et al. (2016); Qin et al. (2023). In Ref. Özel (2006) and Ref. Hoyos et al. (2016), the theoretical modeling of bursting neutron-star spectra and top-down holographic model for strongly interacting quark matter were employed, respectively, to demonstrate that the 2 $M_{\odot}$ neutron star has ruled out soft EOSs of neutron-star matter, and no quark matter exists in massive neutron stars. Recently, it has been argued that as the density increases, the boundaries of hadrons disappear gradually and the corresponding phase transition is a crossover Baym et al. (2018). According to this assumption, the three-window modeling Masuda et al. (2013a, b) in the crossover region was proposed. Many studies has constructed hybrid EOSs in this scheme, and the corresponding maximum masses of hybrid stars are compatible with 2 $M_{\odot}$ojo et al. (2015); Li et al. (2017, 2018a, 2018b, 2022b); Qin et al. (2023).
@@
In addition to theoretical studies of hadron-quark phase transitions and hybrid EOSs, astronomical observations of neutron star masses, radii, and tidal deformability have also placed constraints on numerous EOSs. Some massive neutron star observations such as PSR J0348+0432 Antoniadis et al. (2013) and PSR J0740+6620 Cromartie et al. (2020) require EOSs should not be too soft, but the tidal deformability constrained in BNS merger event GW170817 indicates the EOSs should not be too stiff Abbott and et al. (2017a, 2018). Recently, the joint $M-R$ observations of neutron stars from NASA’s Neutron Star Interior Composition Explorer (NICER) missions have also imposed some constraints on these EOSs Riley et al. (2019); Miller et al. (2019a); Riley et al. (2021); Miller et al. (2021). In Refs. Bauswein et al. (2019); Miao et al. (2020), the authors claim that the gravitational-wave (GW) emission of GW170817 supports a first-order hadron-quark phase transition at supranuclear densities.
@@
It is known that the Bayesian analysis is a good approach to constrain the EOSs. Researchers have obtained important information about the EOS of QCD in this way Ayriyan et al. (2021); Blaschke et al. (2020); Alvarez-Castillo et al. (2016); Pfaff et al. (2022); Alvarez-Castillo et al. (2020); Ayriyan et al. (2019, 2015); Miller et al. (2019b). However, considering that the lagrangian of the NJL model is convenient for numerical calculation, in this work we have performed calculations focusing on the EOS to get the corresponding hybrid star $M-R$ relations and tidal deformability, and then compared them with the relevant neutron star astronomical observations. The model parameters and properties of hybrid stars are constrained as well.
This paper is organized as follows. In Sec.II, the modified NJL model for nonstrange quark matter is briefly introduced, and the EOSs of quark matter are derived. In Sec.III, the Maxwell construction is used to get hybrid EOSs. With recent astronomical observations of neutron star mass, radius, and tidal deformability, we constrain the hybrid EOSs. For comparison, the $M-R$ relations and tidal deformability results of hybrid stars from six representative hybrid EOSs are presented. Finally, a brief summary is given in Sec.IV.
@@ III Structure of hybrid stars
Once the EOS is determined, we can solve the Tolman-Oppenheimer-Volkoff (TOV) equation numerically to get the $M-R$ and mass-central energy density ($M-\epsilon_{\rm c}$) relations. In Fig. 8 and Fig. 9, to study the influence of $G_{1}$ and $B^{1/4}$ on the $M-R$ and $M-\epsilon_{\rm c}$ relations, the corresponding results for (a) $B^{1/4}=153$ MeV and (b) $G_{1}=1.935$ GeV-2 are shown, respectively. In Fig. 8, the most massive quark star with $G_{1}=1.935,1.950,1.972$ GeV-2 ($B^{1/4}=150,153,161$ MeV) is about 1.52 $M_{\odot}$ (1.57 $M_{\odot}$), not satisfying the 2 $M_{\odot}$ constraint. In addition, the quark stars described by the modified NJL model in this work cannot fulfill the recent $M-R$ constraint from the NICER mission (PSR J0740+6620 Miller et al. (2021) and PSR J0030+0451 Miller et al. (2019a)). However, the hybrid EOSs obtained with the Maxwell construction approach in this work can produce hybrid stars in consistent with these astronomical observations, although their quark matter cores are relatively small (about 0.02 $M_{\odot}$).
Figure 8: The $M-R$ relations of quark stars and hybrid stars for (a) $B^{1/4}=153$ MeV and (b) $G_{1}=1.935$ GeV-2. The $M-R$ relation of the hadron star described by BSK21 is shown by the red line. QS1935, QS1950, QS1972 (HS1935, HS1950, HS1972) refer to quark stars (hybrid stars) with $G_{1}=1.935,1.950,1.972$ GeV-2, respectively, and QS150, QS153, QS161 (HS150, HS153, HS161) refer to quark stars (hybrid stars) with $B^{1/4}=150,153,161$ MeV, respectively. The constraints from the NICER mission (PSR J0740+6620 Miller et al. (2021) and PSR J0030+0451 Miller et al. (2019a)) are shown as dark (light) colored areas, referring to the $68\%$ ($95\%$) confidence levels.
@@ IV Summary and discussion
When the current quark mass is chosen as $m=3.5$ MeV, which is estimated by the particle data group, the modification of the coupling constant $G$ in the normal NJL model is helpful, because it is not only consistent with the QCD requirement in essence, but also in agreement with the lattice simulation results of $T_{\rm c}$. In the 2-flavor case when $T_{\rm c}=173\pm 8$ MeV, our calculations indicate that the parameter space of $G_{1}$ is limited to a range of 1.935-1.972 GeV-2, and the corresponding value of $G_{2}$ is from -1.582 to -0.743 GeV-5, implying that the normal 2-flavor NJL model with the four-quark scalar interaction (corresponding to the case of $G_{2}=0$ in our modified NJL model) is inconsistent with the lattice simulation results. For hybrid EOSs, the influence of $G_{1}$ is very small, but the influence of $B^{1/4}$ is obvious.
Considering astronomical observations and the stability of hybrid stars, the parameter $B^{1/4}$ is constrained to be 150-161, 149-158, 148-157 MeV when $G_{1}=1.935,1.950,1.972$ GeV-2, respectively. The quark EOSs constructed with the modified NJL model in this work is soft, and thus cannot satisfy the 2 $M_{\odot}$ constraint of neutron stars and the $M-R$ constraints from NICER missions. It is noted that in some previous studies, the quark matter cores may not exist in compact stars under the Maxwell construction Özel (2006); Hoyos et al. (2016); Qin et al. (2023), or the maximum mass of quark matter cores may be larger than 0.6 $M_{\odot}$ and the BNS in GW170817 can be hybrid stars Ayriyan et al. (2021); Blaschke et al. (2020); Ayriyan et al. (2019); Pfaff et al. (2022). However, with the modified 2-flavor NJL model in this work, the hybrid EOSs with first-order hadron-quark transitions are still in agreement with current neutron star astronomical observations, and pure nonstrange quark matter cores can exist in hybrid stars, possessing a relatively small mass of 0.014-0.026 $M_{\odot}$. According to the hybrid EOSs constrained in this work, the BNS in GW170817 whose masses are estimated to be 1.17-1.36 and 1.36-1.60 $M_{\odot}$ Abbott and et al. (2017a), respectively, may be hadron stars.
@@ Acknowledgements.
This work is supported in part by the national Key Program for Science and Technology Research Development (2023YFB3002500), the National Natural Science Foundation of China (under Grants No. 12005192, No. 12075213, and No. 12233002), the Project funded by China Postdoctoral Science Foundation (No. 2020M672255, No. 2020TQ0287), National SKA Program of China No. 2020SKA0120300, the National Key R&\&&D Program of China (2021YFA0718500), the Natural Science Foundation of Henan Province of China (No. 242300421375), the Natural Science Foundation for Distinguished Young Scholars of Henan Province under grant number 242300421046, the start-up funding from Zhengzhou University. Y.F.H also acknowledges the support from the Xinjiang Tianchi Program.